OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow -y = 3 - x }\) \(\Large \color{MidnightBlue}{\Rightarrow y = -3 + x }\) Get it till now? \(\Large \color{MidnightBlue}{\Rightarrow x - (-3x + x) = 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow x + 3x - x = 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3x = 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 1 }\) Now, \(\Large \color{MidnightBlue}{\Rightarrow 1 - y = 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow -y = 2 }\) \(\Large \color{MidnightBlue}{\Rightarrow y = -2 }\)

OpenStudy (anonymous):

Great